MKTEXTFM(1) MKTEXTFM(1) NAME mktextfm−createaTFMfileforafont SYNOPSIS mktextfm[options]font DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualKpathsea:Alibraryforpathsearching. mktextfmisusedtogenerateatfmfilefromtheMETAFONTsourcefilesforfont,ifpossible.If destdirisgiven,thegeneratedfilewillbeinstalledthere,otherwisea(rathercomplicated)heuristicisused. Thenameofthegeneratedfileisprintedonstandardoutput. mktextfmistypicallycalledbyotherprograms,ratherthanfromthecommandline. OPTIONS mktextfmacceptsthefollowingoptions: --destdirdir Adirectoryname.Ifthedirectoryisabsolute,itisusedas-is.Otherwise,itisappended totherootdestinationdirectorysetinthescript. --helpPrinthelpmessageandexit. --version Printversioninformationandexit. SEEALSO mktexmf(1),mktexpk(1). Kpathsea7.32x 4January1998 1